Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2016-08-01

Observing: Morgan Fraser, Tom Reynolds
Support Team: Seppo Mattila
TAT: Lluís Galbany, Stefan Taubenberger


Weather conditions

Start of night:  Clear, seeing ~1"
Middle of night: Clear, seeing ~1"-1.2"
End of night: Clear, seeing <1"

------------------------------------------------

Calibrations

Evening: SOFI flats
Morning: EFOSC BR flats

------------------------------------------------


Start of night: SOFI had problems with switching to the off field in our first observations, delayed first observations
23:20    SN2016adj (follow-up): JHK imaging, BG & RG spectroscopy. Due to SOFI problems, there are about 10 additional J exposures.
23:36    End of twilight
01:02    Telluric std: HIP68867
01:08    Flux standard: EG274   
01:23    SN2016aiy (follow-up): JHK imaging
02:50    LSQ15adm (follow-up): JHK imaging
04:18    Switched to EFOSC
04:45    LSQ15adm (follow-up): 2x2700s Gr13. Seems to be very little light, couldn't extract. Too faint, so we didn't take a third spectrum
06:30    SN2013ew (follow-up): 1500s Gr11, galaxy spectrum for redshift
07:01    Switched to SOFI
07:14    ASASSN15og (follow-up): JHK imaging
08:40    OGLE15xl (follow-up): JHK imaging
10:08    Switched to EFOSC
            Image analysis
10:12    OGLE15xl (follow-up): Observed 4xi band in twilight

Comments